# Reward-Shifted Speculative Sampling Is An Efficient Test-Time Weak-to-Strong Aligner

### v1

@@ 1 Introduction
To mitigate these issues, test-time alignment (Khanov et al., 2024; Li et al., 2024a; Qiu et al., 2024) has emerged as a training-free alternative to RLHF, allocating more compute during inference to improve alignment quality. Nevertheless, most test-time alignment approaches require interaction between the LLM and an external reward model (RM) throughout decoding, resulting in either excessive LLM calls (best-of- and rejection sampling) or too many RM calls (Deng and Raffel, 2023; Khanov et al., 2024). Although recent studies have recognized these inefficiencies and proposed acceleration techniques (Li et al., 2024a; Qiu et al., 2024; Sun et al., 2024), these methods still rely on external RMs and require numerous calls to large models.
@@ 2.1 RLHF and Reward-Shifted Decoding
Built upon the above objective, another direct approach is to sample multiple responses and only keep the responses that satisfy the reward constraint, known as test-time alignment (Khanov et al., 2024; Li et al., 2024a; Qiu et al., 2024). Test-time alignment only modifies the decoding process of base models to pursue higher reward, namely a reward-shifted decoding process. There are two basic approaches to reward-shifted decoding: i) best-of-, which generates multiple candidates in parallel and selects only the best one: , and ii) rejection sampling, which continues generating proposals until a reward threshold is met: . These approaches are often applied to different granularities (Khanov et al., 2024; Li et al., 2024a) and even combined together (Qiu et al., 2024; Sun et al., 2024) for effective and efficient alignment.
@@ 3 Methodology
Most of test-time alignment methods require a reward model as the signal of human preference (Khanov et al., 2024; Li et al., 2024a; Qiu et al., 2024). The additional computation induced by such a two-model decoding process makes test-time alignment not efficient enough for time-intensive LLM serving (Wang et al., 2024b), which calls for new frameworks to further accelerate the reward-guided decoding process. We are inspired by speculative sampling (Chen et al., 2023), which accelerates autoregressive decoding via a draft-then-verify procedure. We propose to shift the small draft model to align with human preferences (computationally cheap) and design a new speculative sampling algorithm to simulate the distribution of an aligned target model without actually obtaining it. We also prove that the proposed algorithm recovers the RLHF optimal solution.
@@ 3.2 Speculative Sampling under Draft-Target Distributional Shift
The aforementioned distributional shift caused by aligning draft models to human preferences is a critical problem in reward-shifted speculative sampling. To resolve this, we propose a new speculative sampling algorithm that utilizes such a distributional shift to recover the optimal solution of RLHF, which is typically obtainable by training the target model on preference data (Schulman et al., 2017; Rafailov et al., 2023; Shao et al., 2024). We eliminate the need for training target models and significantly reduce the decoding cost compared with previous test-time alignment methods (Khanov et al., 2024; Li et al., 2024a; Qiu et al., 2024).
@@
4 Empirical Results
We provide comprehensive experimental results in Section C, to evaluate the alignment quality and efficiency of the propose method. Our method improves the gold reward at a significantly reduced inference cost compared to other test-time alignment methods.
5 Conclusion and Limitations
Statement of Reproducibility
Statement of AI Assistant Usage
Appendix A Related Work
A.1 Test-Time Alignment
Early alignment research focused on post-training LLMs with techniques like RLHF (Ouyang et al., 2022), but the high cost of policy optimization methods like PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024) has motivated the test-time (or decoding-time) alignment approaches that leave the LLMs’ parameters frozen. For example, best-of- (BoN) generates multiple complete responses and select the highest-reward one; rejection sampling continues generating proposal responses until a reward threshold is met. Built upon these 2 techniques, test-time alignment methods at varying granularities are developed. For example, token-level reward-guided search (Deng and Raffel, 2023; Khanov et al., 2024), segment-level rejection sampling (Li et al., 2024a), and segment-level MC tree search (Qiu et al., 2024) have all been explored. However, although they improve the efficiency of test-time alignment, they still rely on external reward models to interact with the decoding process, which is the major cause of the inefficiency. Our algorithm (SSS) jumps out of the frameworks with reward models, and leverages a small draft model to reflect human preference. The interaction between draft and target models are far more efficient than the traditional LLM-RM framework.
A.2 Speculative Sampling
Orthogonal to test-time alignment, speculative decoding (Chen et al., 2023; Leviathan et al., 2023) accelerates LLM generation by using a lightweight draft model to guess future tokens, which is then verified by a large target model in parallel. Speculative sampling recovers the exact distribution of the target model (Chen et al., 2023). Recent work has aimed to further improve the efficiency of this draft–verify framework. One major direction is increasing the number of candidate tokens accepted by the target model. To this end, tree-based methods (Miao et al., 2024; Li et al., 2024b; Fu et al., 2024) generate multiple draft token paths in parallel, increasing the likelihood of acceptance and reducing verification overhead. Other studies enhance draft token quality through knowledge distillation (Zhou et al., 2023; Liu et al., 2024), layer-skipped decoding (Elhoushi et al., 2024; Zhang et al., 2024), or by adding specialized speculative heads such as MEDUSA (Cai et al., 2024) to improve token prediction and verification efficiency. Some recent variants also explore using reward models (RMs) as verifiers instead of target LLMs (Liao et al., 2025). This change allows for higher token acceptance rates by relaxing fluency constraints. Nakshatri et al. (2024) applies speculative sampling to accelerate constrained decoding, still relying on external reward models to enforce the constraints. Despite these advances, existing speculative decoding methods are primarily designed for computational acceleration and assume an unaligned draft model. As a result, the question of how to incorporate human preference alignment into the speculative decoding process without sacrificing efficiency remains largely under-explored.
Appendix B Proofs
B.1 RLHF Optimal Solution
B.2 Theorem 1
The left half can be re-written as:
For the right half, we first transform the probability of rejection to be:
Then, the probability of generating after draft token rejection is:
always equivalent to the RLHF optimal solution.
Appendix C Experiments
In this section, we present the experimental settings and empirical results. We first discuss the configurations of experiments in Section C.1, then show the main empirical results in Section C.2, and finally have ablation studies on the draft model selection and alternative algorithm options in Section C.3.
C.1 Experimental Settings
@@
For inference experiments, we use the Transformers library222https://github.com/huggingface/transformers. with a temperature of 0.8 and a maximum sequence length of 128 tokens. These settings are standard in previous works (Chen et al., 2023; Khanov et al., 2024; Li et al., 2024a).
Additionally, due to the small size of draft models, all experiments can be conducted on one NVIDIA L40S GPU333https://www.nvidia.com/en-us/data-center/l40s.. Time measurements are obtained from the time.time() API in Python444https://docs.python.org/3/library/time.html..
C.2 Alignment Quality and Inference Efficiency
Table˜4 demonstrates the performance advantages of our method (SSS) across different model pairs on HH-RLHF for a single run. We compare against a range of test-time alignment baselines, including Best-of- (BoN), TreeBoN (Qiu et al., 2024), and CARDS (Li et al., 2024a). We keep the reward model sizes used in these baselines the same as draft models for fair comparison. The baseline choices represent a spectrum of test-time alignment strategies balancing alignment quality and computational efficiency. We also include comparisons with a standard speculative sampling (SS) approach.
@@ C.2 Alignment Quality and Inference Efficiency
C.2.1 Why SSS achieves alignment quality and inference efficiency simutenously?
@@
Draft model alignment reduces reward evaluation cost. By aligning a small draft model to human preference using DPO (Rafailov et al., 2023), SSS avoids frequent reward model queries during decoding, significantly lowering computational overhead compared to segment-level or prefix-level alignment methods like CARDS (Li et al., 2024a) and TreeBoN (Qiu et al., 2024).
@@
C.3 Ablation Studies
C.3.1 How to select a good draft model?
@@
C.3.2 Practical tricks to handle the gap between DPO draft models and the desired well-aligned draft model

### v2

@@ 1 Introduction
To mitigate these issues, test-time alignment (Khanov et al., 2024; Li et al., 2025; Qiu et al., 2024) has emerged as a training-free alternative to RLHF, allocating more compute during inference to improve alignment quality. Nevertheless, most test-time alignment approaches require interaction between the LLM and an external reward model (RM) throughout decoding, resulting in either excessive LLM calls (best-of- and rejection sampling) or too many RM calls (Deng and Raffel, 2023; Khanov et al., 2024). Although recent studies have recognized these inefficiencies and proposed acceleration techniques (Li et al., 2025; Qiu et al., 2024; Sun et al., 2024), these methods still rely on external RMs and require numerous calls to large models.
@@ 2.1 RLHF and Reward-Shifted Decoding
Built upon the above objective, another direct approach is to sample multiple responses and only keep the responses that satisfy the reward constraint, known as test-time alignment (Khanov et al., 2024; Li et al., 2025; Qiu et al., 2024). Test-time alignment only modifies the decoding process of base models to pursue higher reward, namely a reward-shifted decoding process. There are two basic approaches to reward-shifted decoding: i) best-of-, which generates multiple candidates in parallel and selects only the best one: , and ii) rejection sampling, which continues generating proposals until a reward threshold is met: . These approaches are often applied to different granularities (Khanov et al., 2024; Li et al., 2025) and even combined together (Qiu et al., 2024; Sun et al., 2024) for effective and efficient alignment.
@@ 3 Methodology
Most of test-time alignment methods require a reward model as the signal of human preference (Khanov et al., 2024; Li et al., 2025; Qiu et al., 2024). The additional computation induced by such a two-model decoding process makes test-time alignment not efficient enough for time-intensive LLM serving (Wang et al., 2024b), which calls for new frameworks to further accelerate the reward-guided decoding process. We are inspired by speculative sampling (Chen et al., 2023), which accelerates autoregressive decoding via a draft-then-verify procedure. We propose to shift the small draft model to align with human preferences (computationally cheap) and design a new speculative sampling algorithm to simulate the distribution of an aligned target model without actually obtaining it. We also prove that the proposed algorithm recovers the RLHF optimal solution.
@@ 3.2 Speculative Sampling under Draft-Target Distributional Shift
The aforementioned distributional shift caused by aligning draft models to human preferences is a critical problem in reward-shifted speculative sampling. To resolve this, we propose a new speculative sampling algorithm that utilizes such a distributional shift to recover the optimal solution of RLHF, which is typically obtainable by training the target model on preference data (Schulman et al., 2017; Rafailov et al., 2023; Shao et al., 2024). We eliminate the need for training target models and significantly reduce the decoding cost compared with previous test-time alignment methods (Khanov et al., 2024; Li et al., 2025; Qiu et al., 2024).
@@
4 Experiments
In this section, we present the experimental settings and empirical results. We first discuss the configurations of experiments in Section 4.1, then show the main empirical results in Section 4.2, and finally have ablation studies on the draft model selection and alternative algorithm options in Section D.
4.1 Experimental Settings
@@
For inference experiments, we use the Transformers library222https://github.com/huggingface/transformers. with a temperature of 0.8 and a maximum sequence length of 128 tokens. These settings are standard in previous works (Chen et al., 2023; Khanov et al., 2024; Li et al., 2025).
Additionally, due to the small size of draft models, all experiments can be conducted on one NVIDIA L40S GPU. Time measurements are obtained from the time.time() API in Python.
4.2 Alignment Quality and Inference Efficiency
Table˜4 demonstrates the performance advantages of SSS across different model pairs on HH-RLHF for a single run. We compare against a range of test-time alignment baselines, including Best-of- (BoN), TreeBoN (Qiu et al., 2024), and CARDS (Li et al., 2025). We keep the reward model sizes used in these baselines the same as draft models for fair comparison. The baseline choices represent a spectrum of test-time alignment strategies balancing alignment quality and computational efficiency. We also include comparisons with a standard speculative sampling (SS) approach.
@@ 4.2 Alignment Quality and Inference Efficiency
5 Conclusion and Limitations
Statement of Reproducibility
Statement of AI Assistant Usage
Appendix A Related Work
A.1 Test-Time Alignment
Early alignment research focused on post-training LLMs with techniques like RLHF (Ouyang et al., 2022), but the high cost of policy optimization methods like PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024) has motivated the test-time (or decoding-time) alignment approaches that leave the LLMs’ parameters frozen. For example, best-of- (BoN) generates multiple complete responses and select the highest-reward one; rejection sampling continues generating proposal responses until a reward threshold is met. Built upon these 2 techniques, test-time alignment methods at varying granularities are developed. For example, token-level reward-guided search (Deng and Raffel, 2023; Khanov et al., 2024), segment-level rejection sampling (Li et al., 2025), and segment-level MC tree search (Qiu et al., 2024) have all been explored. However, although they improve the efficiency of test-time alignment, they still rely on external reward models to interact with the decoding process, which is the major cause of the inefficiency. Our algorithm (SSS) jumps out of the frameworks with reward models, and leverages a small draft model to reflect human preference. The interaction between draft and target models are far more efficient than the traditional LLM-RM framework.
A.2 Speculative Sampling
Orthogonal to test-time alignment, speculative decoding (Chen et al., 2023; Leviathan et al., 2023) accelerates LLM generation by using a lightweight draft model to guess future tokens, which is then verified by a large target model in parallel. Speculative sampling recovers the exact distribution of the target model (Chen et al., 2023). Recent work has aimed to further improve the efficiency of this draft–verify framework. One major direction is increasing the number of candidate tokens accepted by the target model. To this end, tree-based methods (Miao et al., 2024; Li et al., 2024; Fu et al., 2025) generate multiple draft token paths in parallel, increasing the likelihood of acceptance and reducing verification overhead. Other studies enhance draft token quality through knowledge distillation (Liu et al., 2024), layer-skipped decoding (Elhoushi et al., 2024; Zhang et al., 2024), or by adding specialized speculative heads such as MEDUSA (Cai et al., 2024) to improve token prediction and verification efficiency. Some recent variants also explore using reward models (RMs) as verifiers instead of target LLMs (Liao et al., 2025). This change allows for higher token acceptance rates by relaxing fluency constraints. Nakshatri et al. (2025) applies speculative sampling to accelerate constrained decoding, still relying on external reward models to enforce the constraints. Despite these advances, existing speculative decoding methods are primarily designed for computational acceleration and assume an unaligned draft model. As a result, the question of how to incorporate human preference alignment into the speculative decoding process without sacrificing efficiency remains largely under-explored.
Appendix B Proofs
B.1 RLHF Optimal Solution
B.2 Theorem 1
The left half can be re-written as:
Here, (a) is from Eq. (6), where we design the specific form for the proof completeness. For the right half, we first transform the probability of rejection to be:
Then, the probability of generating after draft token rejection is:
always equivalent to the RLHF optimal solution.
Appendix C Why SSS Achieves Alignment Quality and Inference Efficiency Simultaneously?
@@
Draft model alignment reduces reward evaluation cost. By aligning a small draft model to human preference using DPO (Rafailov et al., 2023), SSS avoids frequent reward model queries during decoding, significantly lowering computational overhead compared to segment-level or prefix-level alignment methods like CARDS (Li et al., 2025) and TreeBoN (Qiu et al., 2024).
@@
Appendix D Ablation Studies
D.1 How to select a good draft model?
@@
D.2 Practical tricks to handle the gap between DPO draft models and the desired well-aligned draft model
@@
D.3 Is SSS Really Approaching the RLHF Optimal Solution?
As this paper claims that SSS can recover the RLHF optimal solution, we compare two RLHF-trained target models (DPO333https://huggingface.co/Nagi-ovo/Llama-3-8B-DPO. and PPO444https://huggingface.co/OpenRLHF/Llama-3-8b-rlhf-100k.) with SSS in Table 6. The results demonstrate that SSS achieves a better gold reward than the trained target models. It is noteworthy that the RLHF-trained target models are all approximations to the optimal solution, and therefore are they presenting lower gold rewards than SSS.
| Method | Gold Reward |
| DPO | 4.85 |
| PPO | 5.62 |
| SSS (ours) | 6.14 |